A new type of rope has a mean breaking strength of 15 kilograms with a standard deviation of 0.5 kilogram. To test the hypothesis, a random sample of 50 lines are tested and the average breaking strength is found to be 14.8 kilograms.

(a) Test the hypothesis, at the 0.01 level of significane the the mean is that μ = 15 kilograms against the alternative that μ < 15 kilograms.

(b) Evaluate the P value of the test.
Answer #1

Answer:

//img.homeworklib.com/questions/6a2d40c0-fd4a-11ea-af96-e5d88fc3e492.png?x-oss-process=image/resize,w_560= 15,  //img.homeworklib.com/questions/6be269d0-fd4a-11ea-a41a-43e434b7aa20.png?x-oss-process=image/resize,w_560 = 0.5

n= 50, T = 14.8

//img.homeworklib.com/questions/6f143e90-fd4a-11ea-a4f5-3bac5e379c58.png?x-oss-process=image/resize,w_560= 0.01

null and alternative hypothesis is

Ho:  1596119865863_blob.png = 15

H1:  1596119865904_blob.png < 15

formula for test statistics is

1596119865943_blob.png

z = \frac{14.8 - 15 }{ 0.5 / \sqrt{50}}

z = −2.828

test statistics : Z = - 2.83

calculate the z critical value for left tailed test with //img.homeworklib.com/questions/6f143e90-fd4a-11ea-a4f5-3bac5e379c58.png?x-oss-process=image/resize,w_560= 0.01

Critical value = -2.33

Decision rule is

Reject Ho if ( test statistics ) 1596119865959_blob.png ( -2.33 )

here, ( -2.83 ) 1596119865913_blob.png ( -2.33 )

Hence, Null hypothesis is rejected.

There is enough evidence to conclude that the population mean is less than 15 kilogram.

b)

calculate P-Value

P-Value = P(z < - 2.83 )

using normal z table we get

P(z < -2.83) = 0.0023

P-Value = 0.0023

P-Value = 0.0023